OpenStudy (anonymous):

The value of 5/(9x^2 + 6x + 11) is Minimum when x=?

OpenStudy (anonymous):

The value of 5/(9x^2 + 6x + 11) is Maximum when x=? --------

OpenStudy (anonymous):

Minimum

OpenStudy (anonymous):

for max we will diff the deno

hartnn (hartnn):

when denom=9x^2+6x+11 is max........so 18x+6=0 implies x=-1/3

OpenStudy (anonymous):

for min wat will we do

OpenStudy (anonymous):

(9x^2 + 6x + 11) find the maximum value of this... use completed square form

OpenStudy (anonymous):

I think @mukushla is very correct....... because the minimum value will be 0 when x--->infinity

OpenStudy (anonymous):

i just wanted to make it little uglier...lol